Y=C+IO+GO; C=CO+α(Y-T)
T=TO+ty co=100, Io=90 GO=330 TO=240 α=75% t=20% find national income and c
1
Expert's answer
2016-02-17T00:00:57-0500
Y=C+IO+GO; C=CO+α(Y-T)
T=TO+tY
Co=100,
Io=90
GO=330
TO=240
α=75%
t=20%.
National income Y=C + IO + GO = CO + α(Y-T) + IO + GO = CO + α*Y - α*(To + t*Y) + IO + GO = 100 + 0.75Y - 0.75*(240 + 0.2Y) + 90 + 330 = 340 + 0.75 Y - 0.15Y = 340 + 0.6Y, then
0.4Y = 340
Y = 850
c = C/Y
C = Co + α(Y - To - t*Y) = 100 + 0.75(850 - 240 - 0.2*850) = 430
c = 430/850 = 0.51